OpenStudy (anonymous):

I'm having a problem with algebra. I'm supposed to simplify this problem: 3x^2 + 7x +2 OVER 4x^2-16. Anyone know how to start? I'm trying to break down the top into theFOIL method but having a hard time figuring out what 2 numbers are going to get me a positive 7x and a positive 2.

OpenStudy (amistre64):

forget that; i lost my place in reading

OpenStudy (amistre64):

3x^2 + 7x +2 first times last 3*2 = ?

OpenStudy (anonymous):

6

OpenStudy (amistre64):

good; the last sign is + which tells us we add; the middle sign is + which tells us we add + and + (x+ ?/3) (x+ ?/3) what are the factors of 6 that add up to 7?

OpenStudy (anonymous):

6 and 1

OpenStudy (amistre64):

good, then lets fill; them in (x+ 6/3) (x+ 1/3) reduce the fractions (x+ 2) (x+ 1/3) if theres still a fraction stick the bottom in front (x+ 2) (3x+ 1) and we are done with it

OpenStudy (amistre64):

factor the bottom part and cancel out like "terms"

OpenStudy (anonymous):

I don't think I can simplify the bottom anymore. It is 4x^2-16. I have it simplified as (2x-4)(2x+4) So I think my answer would just be (x+2)(3x+1) OVER 2x-4)(2x+4). Id that right?

OpenStudy (amistre64):

factor out a 2 from each one

OpenStudy (amistre64):

ideally youd factor out a 4 first; but either way you get to it in the end :)

OpenStudy (anonymous):

Ok. Totally lost now. Sorry.

OpenStudy (amistre64):

can you factor out a 2 from each of your denominator parts?

OpenStudy (amistre64):

4x^2 - 16 = 4 (x^2 - 4) = 4 (x-2) (x+2) or your way 4x^2 - 16 = (2x-4) (2x+4) = 2(x-2) 2(x+2) = 4 (x-2) (x+2)

OpenStudy (anonymous):

I see it. So then I could cancel out the x+2 from the top and bototm.

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Got it! Even more important, I understand it!! Thank you!